Determine the [H+] , [OH−] , and pH of a solution with a pOH of 2.81 at 25 °C.

[H+]= 

[OH−]= 

pH= 



1
Expert's answer
2022-04-26T05:07:03-0400

"[H^+]= 10^{-pH}=10^{-(14-pOH)}=10^{-11.19}=6.46*10^{-12}M"

"[OH^-]= 10^{-pOH}=10^{-2.81}=1.55*10^{-3}M"

"pH=14-pOH=11.19"